# Krylov complexity as an order parameter for quantum chaotic-integrable transitions

### v2

@@ Abstract
Discussion. In this Letter, we proposed that the KCP may serve as a defining feature of quantum chaos and could act as a robust ‘order parameter’ for identifying quantum chaotic phases in many-body systems. We observed that the KCP vanishes in integrable systems while displaying critical dynamics across chaotic-to-integrable quantum and thermal transitions. By computing the KCP using the TFD as an initial state, we effectively identified the chaotic-integrable transitions in the mass-deformed SYK model at both infinite and finite temperature. This finding is consistent with the results obtained from traditional probes such as spectral statistics and out-of-time-order correlators, and aligns with the independent prediction from the SFF.
@@
Lastly, our analysis underscores the unique importance of the TFD state in the study of quantum chaos in many-body systems, building on previous observations Camargo et al. (2024b). This state is crucial for Krylov complexity to effectively probe random matrix physics, likely because random matrix behavior encompasses the entire level-spacing spectrum of the system. Consequently, any measure aimed at probing random matrix behavior benefits from the TFD state’s ability to encompass the full spectrum. TFD states also play a significant role in holography, serving as CFT duals of two-sided black holes Maldacena (2003). Recent findings suggest a connection between the Krylov complexity of chord states in the double-scaled SYK model and the length of the dual Lorentzian wormhole in Jackiw-Teitelboim gravity Rabinovici et al. (2023), a lower-dimensional realization of holography. Therefore, our work might provide valuable insights into the holographic dual description of chaotic-integrable transitions in quantum systems and contribute to addressing the profound and long-standing ‘million-dollar question’ of quantum gravity.
Note added: While this paper was in preparation, Bhattacharjee and Nandy (2024); Chapman et al. (2024); Scialchi et al. (2024b) appeared, partially overlapping with our scope.
@@ Acknowledgements.
Supplemental material
In this Supplementary Material, we offer additional details on the numerical computations discussed in the main text and present further analyses to corroborate our findings.
I Block diagonalization of the SYK Hamiltonian
@@
II Lanczos coefficients and quantum chaos
@@
III Normalization condition and Ehrenfest theorem
@@
Figure 8: The normalization condition of the Krylov wave functions.
@@
$\begin{split}\partial^{2}_{t}C(t)=-\left[\left[C(t)\,,\mathcal{L}\right],\mathcal{L}\right]\,,\end{split}$ (16)
where $\mathcal{L}=H\otimes\mathbb{I}$ denotes the Liouvillian and $\mathbb{I}$ represents the identity operator. By applying the Schrödinger equation along with the definition of Krylov complexity, the Ehrenfest theorem (16) can be formulated in terms of the Lanczos coefficients and the Krylov wave functions as follows:
$\begin{split}\partial^{2}_{t}C(t)=2\sum_{n}\left[\left(b_{n+1}^{2}-b_{n}^{2}\right)\psi_{n}(t)\psi_{n}^{*}(t)+\left(a_{n+1}-a_{n}\right)b_{n+1}\psi_{(n+1}(t)\psi_{n)}^{*}(t)\right]\,,\end{split}$ (17)

### v4

@@ V Discussion
In this paper, we proposed that the KCP may serve as a defining feature of quantum chaos and could act as a robust ‘order parameter’ for identifying quantum chaotic phases in many-body systems. We observed that the KCP vanishes in integrable systems while displaying critical dynamics across chaotic-to-integrable quantum and thermal transitions. By computing the KCP using the TFD as an initial state, we effectively identified the chaotic-integrable transitions in the mass-deformed SYK and sparse SYK models at both infinite and finite temperature. This finding is consistent with the results obtained from traditional probes such as spectral statistics and out-of-time-order correlators, and aligns with the independent prediction from the SFF.
@@
Lastly, our analysis underscores the importance of the TFD state in the study of quantum chaos in many-body systems, building on previous observations Camargo et al. (2024b). This state is crucial for Krylov complexity to effectively probe random matrix physics, likely because random matrix behavior encompasses the entire level-spacing spectrum of the system. Consequently, any measure aimed at probing random matrix behavior benefits from the TFD state’s ability to encompass the full spectrum. TFD states also play a significant role in holography, serving as CFT duals of two-sided black holes Maldacena (2003). Recent findings suggest a connection between the Krylov complexity of chord states in the double-scaled SYK model and the length of the dual Lorentzian wormhole in Jackiw-Teitelboim gravity Rabinovici et al. (2023), a lower-dimensional realization of holography. Therefore, our work might provide valuable insights into the holographic dual description of chaotic-integrable transitions in quantum systems and contribute to addressing the profound and long-standing ‘million-dollar question’ of quantum gravity.
Furthermore, it may be worthwhile to explore the connection between Krylov complexity and more classical approaches to integrability. One interesting direction could involve an “extended” definition of quantum chaos inspired by the Liouville-Arnold theorem, stating that a system with $N$ degrees of freedom is considered regular if the number of the first integrals, $M$, equals $N$, while the system is deemed chaotic when $M$ is less than $N$.
@@ Acknowledgements.
Appendix A Details on the numerical methods
In this appendix, we provide additional details on the numerical computations discussed in the main text, along with further analyses to substantiate our results. Here, we primarily focus on the mass-deformed SYK model; however, the key features are consistent with those observed in the sparse SYK model as well.
A.1 Block diagonalization of the SYK Hamiltonian
@@
A.2 Lanczos coefficients and quantum chaos
@@
A.3 Normalization condition and Ehrenfest theorem
@@
Figure 10: The normalization of the Krylov wave functions.
@@
$\begin{split}\partial^{2}_{t}\langle\psi|\hat{C}|\psi\rangle=-\langle\psi|\left[\left[\hat{C}\,,\mathcal{L}\right],\mathcal{L}\right]|\psi\rangle\,\end{split}$ (19)
where $\hat{C}:=\sum_{n}n\,|K_{n}\rangle\langle K_{n}|$ and $\mathcal{L}=H\otimes\mathbb{I}$ denotes the Liouvillian and $\mathbb{I}$ represents the identity operator. By applying the Schrödinger equation along with the definition of Krylov complexity, the Ehrenfest theorem (19) can be formulated in terms of the Lanczos coefficients and the Krylov wave functions as follows:
$\begin{split}\partial^{2}_{t}C(t)&=2\sum_{n}\left[\left(b_{n+1}^{2}-b_{n}^{2}\right)\psi_{n}(t)\psi_{n}^{*}(t)\right.\\ &\left.+\left(a_{n+1}-a_{n}\right)b_{n+1}\psi_{(n+1}(t)\psi_{n)}^{*}(t)\right]\,,\end{split}$ (20)
@@
[…]
A.4 The depth of the SFF-hole
We further explore the dynamics of the dip in the SFF as the system transitions from chaotic to integrable regimes. Specifically, we observe that the depth of the dip tends to be suppressed along this transition.
To quantify this behavior, we utilize Eq. (9) and define
$\begin{split}\Delta\text{SFF}=\text{Log}\left[\text{SFF}(t_{\text{dip}})\right]-\text{Log}\left[\text{SFF}(t=\infty)\right]\,,\end{split}$ (21)
which characterizes the difference between the dip and the saturation value of the SFF.
Fig. 12 presents the computed $\Delta\text{SFF}$ for the SYK models discussed in the main text. A comparison with the KCP, as shown in the bottom panels of Figs. 1 and 4, reveals a notable similarity in the behavior. However, we find that the KCP serves as a more robust indicator of the chaotic-to-integrable transition. In particular, $\Delta\text{SFF}$ fails to effectively capture the transition in the limit of $\beta=0$ for the mass-deformed SYK model (red data in the top panel of Fig. 12). This is not a surprise as the depth of the dip is still manifest for the SYK2 model (corresponding to $\kappa\rightarrow\infty$ case) in the $\beta=0$ Erdmenger et al. (2023a).